# An SEIR network epidemic model with manual and digital contact tracing allowing delays

### v2

@@ 4.2 The effectiveness of manual and digital contact tracing
If manual and digital contact tracing would have acted independently, then the reproduction number for the combined model would have been $R_{0}(1-r_{M})(1-r_{D}).$ This raises a question: How does $(1-r_{DM})$ compare to the product $(1-r_{M})(1-r_{D})$?
@@
Figure 4.2: a): Plot of critical combination $(p_{M},\pi_{A})$ such that $R_{MD}=1$ (solid curves) and $R_{0}(1-r_{M})(1-r_{D})=1$ (dashed curves), where $p_{D}=0.8$ and $T_{D}\equiv 3$ are fixed; and the contact rates $\beta_{L},\beta_{G}$ are chosen for $\alpha=50\%,75\%$; $R_{0}=3$ and $1.5$. b): Plot of the reductions $r_{M},r_{D}$ and $r_{MD}$ of $R_{0}$ against the fraction $\alpha$ of network transmission $\in[0.01,1]$, where $R_{0}=3,$ $p_{D}=0.8$, $T_{D}\equiv 3$ and $p_{M}=\pi_{A}=0.5$.
@@ 4.3 The effectiveness of manual contact tracing
To conclude, manual contact tracing is only effective if a large fraction of contacts are traced and the tracing delay is short, but it cannot alone reduce $R_{M}$ below 1 if $R_{0}$ is larger than 2 say.
@@ 4.4 The effectiveness of digital contact tracing
A disadvantage with digital contact tracing is that it only happens for contacts where both the infector and infectee are app-users, something which happens with probability $\pi_{A}^{2}$. This is why manual tracing is more effective for small values of $p_{M}=\pi_{A}$, and in most Western countries, the app-using fraction never became very big. In the figure, we also plot the reduction $r_{D}$ as a function of $\pi_{A}^{2}$, which is seen to outperform manual tracing for all values of $\pi_{A}^{2}=p_{M}$.
Figure 4.5: Plot of reduction $r_{M}$ and $r_{D}$ with $p_{M}$, $\pi_{A}$, and $\pi_{A}^{2}$, where $\beta_{L},\beta_{G}$ are chosen so that there are $50\%$ transmission on network when $R_{0}=3$, $p_{D}=0.8$ and $T_{D}\equiv 3$.
@@ 5 Discussion
Despite these insights, several challenges remain to be solved for future work. A limitation of our model is that we restricted the infectious period to be deterministic so we could have a branching process approximation for the early epidemic, which seems harder in general. Further, we have only assumed one step instead of iterative tracing, while in most empirical cases, traced individuals that are infectious would be further questioned, thus possibly triggering manual and/or digital tracing, and so on. Moreover, we have not considered that app-using might have assortative nature, i.e., app-users’ friends are more likely to use the app as well. It would be interesting to see how that influences the overall effectiveness of digital contact tracing. We also ignored the presence of clustering in the social networks, which would have consequences for the impact of both disease spreading and contact tracing [kiss2005cluster, keeling_networks_2005]. An interesting problem not considered in the present paper is fatigue against contact tracing. An empirically observed problem with contact apps was that when incidence grew large, app-users would receive warnings very frequently, thus leading to fewer people following the instructions to self-isolate and test each time. It would be an interesting study to investigate how many tests individuals would be required to do for both types of contact tracing. We emphasize however that the current analysis focuses on the beginning of an epidemic outbreak, and then incidence is low, implying that not many tests will be performed.
Contrasting our model’s pessimistic assumptions with the over-optimistic scenarios of full, iterative and instantaneous contact tracing studied in [zhang2022] indicates that real-world outcomes may lie somewhere in between. Moreover, our findings suggest that the actual combined effect of both manual and digital tracing is, unfortunately, smaller than the product of their separate effects. In conclusion, this paper contributes valuable perspectives on the interplay between manual and digital contact tracing in epidemic control and their respective contributions.

### v4

@@ 4.2 The effectiveness of manual and digital contact tracing
If manual and digital contact tracing would have acted independently, then the reproduction number for the combined model would have been $R_{0}(1-r_{M})(1-r_{D}).$ This raises a question: How does $1-r_{DM}$ compare to the product $(1-r_{M})(1-r_{D})$?
@@
Figure 4.2: a): Plot of critical combination $(p_{M},\pi_{A})$ such that $R_{MD}=1$ (solid curves) and $R_{0}(1-r_{M})(1-r_{D})=1$ (dashed curves), where $p_{D}=0.8$ and $T_{D}\equiv 3$ are fixed, and the contact rates $\beta_{L},\beta_{G}$ are chosen for $\alpha=50\%,75\%$ and $R_{0}=3,1.5$; b): Plot of the reductions $r_{M},r_{D}$ and $r_{MD}$ of $R_{0}$ against the fraction of network transmission $\alpha\in[0.01,1]$, where $R_{0}=3,$ $p_{D}=0.8$, $T_{D}\equiv 3$ and $p_{M}=\pi_{A}=0.5$.
@@ 4.3 The effectiveness of manual contact tracing
To conclude, manual contact tracing is only effective if a large fraction of contacts are traced and the tracing delay is short, but it cannot alone reduce $R_{M}$ below 1 if $R_{0}$ is larger than e.g. 2.
@@ 4.4 The effectiveness of digital contact tracing
A disadvantage with digital contact tracing is that it only happens for contacts where both the infector and infectee are app-users, something which happens with probability $\pi_{A}^{2}$. This is why manual tracing is more effective for small values of $p_{M}=\pi_{A}$, and in most Western countries, the app-using fraction never became very big [watch2020covid]. In the figure, we also plot the reduction $r_{D}$ as a function of $\pi_{A}^{2}$, which is seen to outperform manual tracing for all values of $\pi_{A}^{2}=p_{M}$.
Figure 4.5: Plot of the reduction $r_{M}$ and $r_{D}$ in the reproduction number with the effectiveness of contact tracing represented by $p_{M}$, $\pi_{A}$, and $\pi_{A}^{2}$, where $\beta_{L},\beta_{G}$ are chosen so that there are $50\%$ transmission on network when $R_{0}=3$, $p_{D}=0.8$ and $T_{D}\equiv 3$.
@@ 5 Discussion
Despite these insights, several challenges remain to be solved for future work. A limitation of our model is that we restricted the infectious period to be deterministic so we could have a branching process approximation for the early epidemic, which seems harder in general. It would be interesting to extend the model to consider scenario where the infectious period is random (assuming the same mean). We hypothesize that this would increase the efficacy of contact tracing. This is because individuals with longer infectious periods could potentially be traced well before the end of their infectious period, thereby significantly reducing the effect of such super-spreaders, but to prove this seems non-trivial. Further, we have only assumed one step instead of iterative tracing, while in most empirical cases, traced individuals that are infectious would be further questioned, thus possibly triggering manual and/or digital tracing, and so on. Moreover, we have not considered that app-using might have assortative nature, i.e., app-users’ friends are more likely to use the app as well. It would be interesting to see how that influences the overall effectiveness of digital contact tracing. We also ignored the presence of clustering in the social networks, which would have consequences for the impact of both disease spreading and contact tracing [kiss2005cluster, keeling_networks_2005]. Moreover, our current model does not include contact tracing within household settings. Household structures could potentially enhance the manual contact tracing process, since contacts within a household are more easily identified. An interesting problem not considered in the present paper is fatigue against contact tracing. An empirically observed problem with contact apps was that when incidence grew large, app-users would receive warnings very frequently, thus leading to fewer people following the instructions to self-isolate and test each time. It would be an interesting study to investigate how many tests individuals would be required to do for both types of contact tracing. We emphasize however that the current analysis focuses on the beginning of an epidemic outbreak, when incidence is low, implying that not many tests will be performed. In this paper, we focus mainly on contact tracing. There are other preventive measures that can effectively reduce the reproduction number, some of which can be considered within our model. For example, we can include social distancing by reducing the contact rates $\beta_{L}$ and $\beta_{G}$, leading to a smaller $R_{0}$. Additionally, implementing mass testing could increase the diagnosis probability $p_{D}$, thereby enhancing the effectiveness of contact tracing.
Contrasting our model’s “pessimistic” assumptions (forward only and non-iterative contact tracing) with the over-optimistic scenarios of full, iterative and instantaneous contact tracing studied in [zhang2022] indicates that real-world outcomes may lie somewhere in between. Moreover, our findings suggest that at the early stage of the epidemic, the actual combined effect of both manual and digital tracing is, unfortunately, smaller than the product of their separate effects. In conclusion, this paper contributes valuable perspectives on the interplay between manual and digital contact tracing in epidemic control and their respective contributions.